[Brahms-l] Shift report 20050430 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Apr 30 2005 - 16:04:53 EDT
Shifter leader was RP
Shifters were 


runs:

FS 8 deg 608B run 14348 
statistics for tr6 greater than 500K

FS 8 deg 1219B runs 14351, 14353, 14354
statistcs tr2 50K

problem with L1440 fix

pedestal runs done

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Apr 30 16:05:21 2005

This archive was generated by hypermail 2.1.8 : Sat Apr 30 2005 - 16:05:32 EDT